OpenStudy (anonymous):

solve. round to 4th decimals, separate answer with commas if necessary. (x^2)(2x)-(2^x)=0

OpenStudy (anonymous):

Are you sure you have this entered correctly?

OpenStudy (anonymous):

sorry its \[x ^{2}2^{x}-2^{x}=0\]

OpenStudy (anonymous):

I would factor this.. common factor is 2^x. So you would have 2^x(x^2 - 1) = 0 Then you have the difference of two squares so you have 2^x(x + 1)(x - 1) = 0 So every where there is an "x" there is an answer. So x + 1 would give you an answer of -1 ... x - 1 would give you an answer of 1 When does 2^x = 0?

OpenStudy (anonymous):

x would be 0, so -1 and 1 would work